# A Tale of Two Symmetries: Exploring the Loss Landscape of Equivariant Models

### v1

@@ Abstract
Equivariant neural networks have proven to be effective for tasks with known underlying symmetries. However, optimizing equivariant networks can be tricky and best training practices are less established than for standard networks. In particular, recent works have found small training benefits from relaxing equivariance constraints. This raises the question: do equivariance constraints introduce fundamental obstacles to optimization? Or do they simply require different hyperparameter tuning? In this work, we investigate this question through a theoretical analysis of the loss landscape geometry. We focus on networks built using permutation representations, which we can view as a subset of unconstrained MLPs. Importantly, we show that the parameter symmetries of the unconstrained model has nontrivial effects on the loss landscape of the equivariant subspace and under certain conditions can provably prevent learning of the global minima. Further, we empirically demonstrate in such cases, relaxing to an unconstrained MLP can sometimes solve the issue. Interestingly, the weights eventually found via relaxation corresponds to a different choice of group representation in the hidden layer. From this, we draw 3 key takeaways. (1) Viewing any class of networks in the context of larger unconstrained function space can give important insights on loss landscape structure. (2) Within the unconstrained function space, equivariant networks form a complicated union of linear hyperplanes, each associated with a specific choice of internal group representation. (3) Effective relaxation of equivariance may require not only adding nonequivariant degrees of freedom, but also rethinking the fixed choice of group representations in hidden layers.
@@ 1 Introduction
In our work, we seek to address an even more foundational question: Do equivariance constraints themselves create fundamental obstacles to optimization? In order to answer this question, we analyze the loss landscape geometry of equivariant networks built with permutation representations.
@@
  • By viewing a given architecture as a constrained version of a larger function class, one can obtain important insights on loss landscape geometry
  • Within an unconstrained function class, there may be multiple symmetrically related constrained subspaces which represent the same classes of functions
  • To achieve effective relaxation of equivariant models, one may have to rethink our fixed choice of group representation in the hidden layers
@@ 2 Background
Here, we introduce the key concepts necessary for our work. A more in depth overview can be found in Appendix A including an overview of necessary group theory.
@@ 2.1.1 Group representations and irreps
Let $G$ be a group and $V$ a vector space. A group representation $\rho:G\to\mathrm{GL}(V)$ is irreducible if there is no subspace $W\subset V$ other than 00 such that $gW=W$ for all $g\in G$.
@@ 2.1.2 Nonlinearities and permutation representations
The intuition is that for a p-rep, there is always some choice of basis $b_{i}$ such that $\rho(g)b_{i}=b_{\pi(i)}$. In other words, group actions just permute the basis. The trivial representation is an example since there is only one basis and it remains invariant under group action. The regular representation commonly used in group convolutions is another example since we can associate each basis vector to a group element cohen2016group .
@@ 2.2 Parameter symmetries and loss landscapes
Definition 2.7 (Parameter symmetries (informal)).
Let $\mathcal{F}_{\Theta}=\{f_{\theta}:X\to Y:\theta\in\Theta\}$ be a class of functions parameterized by $\theta\in\Theta$. Any homeomorphism $p:\Theta\to\Theta$ is a local parameter symmetry at $\theta$ if $f_{\theta}=f_{p\theta}$ and we denote the group of such transformations as $P_{\theta}$. If $p$ is a local parameter symmetry for all $\theta\in\Theta$, then we say $p$ is a global parameter symmetry and we denote the group of such transformations as $P$.
@@
For us, we usually consider $\Theta$ as our vector space and subgroups of the global parameter symmetries $H\leq P$.
@@
Let $f:M\to\mathbb{R}$ be a smooth function invariant under $H$. Then if $\Sigma$ is a submanifold, then critical points of $f|_{\Sigma}$ are critical points of $f$.
We see that the loss landscape is invariant under $P$ and in particular, invariant under any subgroup of $P$. For example, if we consider the subgroup $H$ corresponding to permuting two specific neurons, the corresponding fixed point subspace $\Theta_{H}$ corresponds to shared input and output weights for those two neurons. This is effectively a network with one fewer neuron. The principle of symmetric criticality then tells us minima of $L|_{\Theta_{H}}$ is a critical point of $L$, reminiscent of previously known results relating minima of smaller networks to critical points of larger ones (fukumizu2000local, ; simsek2021geometry, ; brea2019weight, ). To connect directly back to these prior works, we note that any transformation of the output weights of these two neurons which leaves the sum unchanged is a local symmetry of $\Theta_{H}$. Hence taking the orbit of $\Theta_{H}$ then gives a larger subspace corresponding to the different ways we may embed the smaller network in a larger one.
@@ 3 Loss landscape of equivariant networks
Here we present the main theoretical results and intuition for why they are true. Proofs can be found in Appendix B. We first make some general statements that apply to general linearly constrained models, not just equivariant ones.
@@ 3.1 Hidden symmetries in constrained models
If the hidden fixed point subspace $\tilde{\Theta}_{H}$ is one dimension lower than the constrained subspace, we have an even stronger statement. In this case, $\tilde{\Theta}_{H}$ in some sense splits $\tilde{\Theta}$ into two halves. We cannot traverse between the two halves via gradient descent because once we reach $\tilde{\Theta}_{H}$ we must get stuck. Generically, we expect the global minima to be in one of these halves so if we are initialized in the "bad" half, then we would never reach the global minima. In particular, we show that there must be another minima.
@@
Suppose $\dim(\tilde{\Theta}_{H})=\dim(\tilde{\Theta})-1$. Suppose there is a minima $\theta_{1}$ of $L|_{\tilde{\Theta}}$ such that $\theta_{1}\notin\tilde{\Theta}_{H}$. Further suppose that there exists some $C$ such that for any direction $\hat{r}\in\tilde{\Theta}$ we have $\hat{r}\cdot\nabla L(C\hat{r})>0$. Then if Hessians of all critical points are nondegenerate, there must exist a distinct minima $\theta_{2}\neq\theta_{1}$ of $L|_{\tilde{\Theta}}$.
@@ 3.2 Hidden fixed point spaces of equivariant networks
Let a $m$-neuron block be a transitive permutation representation. Suppose the weights satisfy the corresponding equivariance constraints. If this $m$-neuron block is reducible, then we can replace the transitive permutation representation with a $n$-neuron transitive permutation representation where $n<m$ and $n|m$.
@@
Let a $m$-neuron block and a $m^{\prime}$-neuron block be distinct transitive permutation representations each of which are irreducible. Suppose the $m+m^{\prime}$-neuron block combining these neurons is reducible. Then they must both be the same permutation representation and we can replace the $m+m^{\prime}$-neuron block with a $m$-neuron point.
@@ 4 Experiments
where $\sigma$ is our nonlinearity. For the experiments here, we use ReLU. Results with other activations can be found in Appendix C. All of our experiments can be run on a laptop.
@@
4.1 Permutation symmetry
In this section, we consider permutation symmetry $S_{n}$ and our example uses $S_{3}$. In Appendix C, we also consider spherical signals used in $\mathrm{O}(3)$ and $\mathrm{SO}(3)$ equivariant networks.
@@ 4.2 Loss landscape visualization
For a $\pi_{n}$ block in the hidden layer, we note there is a hidden symmetry corresponding to permuting within the block. The corresponding hidden fixed point subspace happens exactly the weights to these neurons are all the same. If we have a single $\pi_{n}$ block in the input, then the linear map $\theta_{1}I+\theta_{2}(\mathbf{1}\mathbf{1}^{T}-I)$ is in the hidden fixed point subspace when $\theta_{1}=\theta_{2}$. In particular, the degrees of freedom is reduced exactly by 1 so Theorem 3.3 applies and we expect there to be an additional induced minima. Note one can check that additional $\pi_{0}$ representations in the input or hidden layer do not affect this.
Hence, we choose $S_{3}$ and a $5$-dimensional input consisting of one $\pi_{3}$ rep and 2 trivial $\pi_{0}$ reps. In the hidden layer we choose one $\pi_{0}$ rep and $3$ trivial $\pi_{0}$ reps. We randomly generate weights for the teacher network. We then set all other weights of the student network to equal that of the teacher network and only vary the diagonal and off diagonal weights $\theta_{1}$ and $\theta_{2}$ controlling the mapping from $\pi_{3}$ to $\pi_{3}$. Figure 4 shows the resulting loss landscape for a random seed. We clearly see two distinct minima in the loss landscape. In Appendix C, we initialize teacher weights using different random seeds and see similar results.
@@ 4.3 Training, bad initializations, and parameterization
For each value of $\theta_{1},\theta_{2}$, we trained the student network for 100 steps at a learning rate of $10^{-1}$ using gradient descent on the exact loss. This is usually sufficient for the student network to converge. The resulting final loss is shown in Figure 5a. Note that indeed we see two regions, one which converges to a global minima and one which does not. However, the boundary is not a straight line. It turns out this is because gradient descent behavior depends on parameterization and in particular, on the relative scaling of the parameters (kristiadi2023geometry, ). By rescaling our parameterization to $\sqrt{n-1}\theta_{1}I+\theta_{2}(\mathbf{1}\mathbf{1}^{T}-I)$ and then performing training, we obtain Figure 5b. Here, the boundary between the two regions is indeed a straight line and corresponds to the hidden fixed point space. We emphasize, however, that no matter how we rescale our parameters, there will always be a boundary. In Appendix C, we generate teacher weights using different random seeds and see similar results.
@@ 4.4 Relaxation
Figure 7: Training of unconstrained network initialized at bad minima. We break out of the original equivariant subspace into a symmetrically related equivariant subspace. In particular, the "bad" half of the original constrained subspace is closer to the "good" half of the symmetrically related equivariant subspace. (a) Depicts values of various quantities during training of the relaxed model. (b) Projection of the parameters onto a $3$-dimensional subspace. The planes correspond to various symmetrically related equivariant subspaces color coded by good and bad halves containing the good or bad minima respectively.
@@ 5 Conclusion
Appendix A Group theory background

### v2

@@ Abstract
Equivariant neural networks have proven to be effective for tasks with known underlying symmetries. However, optimizing equivariant networks can be tricky and best training practices are less established than for standard networks. In particular, recent works have found small training benefits from relaxing equivariance constraints. This raises the question: do equivariance constraints introduce fundamental obstacles to optimization? Or do they simply require different hyperparameter tuning? In this work, we investigate this question through a theoretical analysis of the loss landscape geometry. We focus on networks built using permutation representations, which we can view as a subset of unconstrained MLPs. Importantly, we show that the parameter symmetries of the unconstrained model has nontrivial effects on the loss landscape of the equivariant subspace and under certain conditions can provably prevent learning of the global minima. Further, we empirically demonstrate in such cases, relaxing to an unconstrained MLP can sometimes solve the issue. Interestingly, the weights eventually found via relaxation corresponds to a different choice of group representation in the hidden layer. From this, we draw 3 key takeaways. (1) By viewing the unconstrained version of an architecture, we can uncover hidden parameter symmetries which were broken by choice of constraint enforcement (2) Hidden symmetries give important insights on loss landscapes and can induce critical points and even minima (3) Hidden symmetry induced minima can sometimes be escaped by constraint relaxation and we observe the network jumps to a different choice of constraint enforcement. Effective equivariance relaxation may require rethinking the fixed choice of group representation in the hidden layers.
@@ 1 Introduction
In our work, we seek to address an even more foundational question: Do equivariance constraints themselves create fundamental obstacles to optimization? In this paper, we show that our choice of constraint enforcement can hide parameter symmetries of the larger unconstrained model class and that these hidden symmetries can induce critical points and even spurious minima.
@@
Figure 1: (a) Standard viewpoint of why constraints may help. If we know the ground truth function satisfies some type of constraint, then enforcing such constraints may reduce the search space allowing for better training. (b) In reality, unconstrained networks contain many parameter symmetries. However, enforcing constraints (such as equivariance) require choices that breaks some of these symmetries. These broken symmetries are "hidden" to the constrained model but still influence gradient values and in some cases create spurious minima and hinder learning. Further, we empirically observe that when completely relaxing equivariance constraints, a network can sometimes "jump" between subspaces corresponding to different ways of enforcing the same type of constraint.
@@
  • By viewing a given architecture as a constrained version of a larger function class, we can uncover previously hidden parameter symmetries that were broken by our choice of constraint enforcement.
  • Hidden symmetries have important consequences on loss landscape and can induce critical points and even minima.
  • Empirically, we find a spurious minimum in an equivariantly constrained subspace can escape to a better minimum in a symmetrically related subspace through constraint relaxation. To achieve more effective relaxation of equivariant models, one may have to rethink our fixed choice of group representation in the hidden layers
@@ 2 Background
Here, we introduce the key concepts necessary for our work. A more in depth overview can be found in Appendix C including an overview of necessary group theory.
@@ 2.1.1 Group representations and irreps
Let $G$ be a group and $V$ a vector space. A group representation $\rho:G\to\mathrm{GL}(V)$ is irreducible if there is no subspace $W\subset V$ other than 0 such that $gW=W$ for all $g\in G$.
@@ 2.1.2 Nonlinearities and permutation representations
Let $G$ be a group. Any group homomorphism $\pi:G\to\mathrm{Sym}(n)$ is an abstract permutation representation. Let $\rho:\mathrm{Sym}(n)\to\mathrm{GL}(V_{n})$ map corresponding permutation matrices acting on $n$-dimensional vector space $V_{n}$. Then $\rho\circ\pi:G\to\mathrm{GL}(V_{n})$ is a permutation representation. We will abbreviate permutation representation as p-rep and $\rho\circ\pi$ to $\rho$.
The intuition is that for a p-rep, there is always some choice of basis $b_{i}$ such that $\rho(g)b_{i}=b_{\pi(i)}$. In other words, group actions just permute the basis. The trivial representation is an example since there is only one basis and it remains invariant under group action. The regular representation commonly used in group convolutions is another example since we can associate each basis vector to a group element cohen2016group . In Appendix C.1, we discuss how to generalize p-reps to infinite dimensions.
@@ 2.2 Parameter symmetries and loss landscapes
Definition 2.7 (Parameter symmetries).
Let $\mathcal{F}_{\Theta}=\{f_{\theta}:X\to Y:\theta\in\Theta\}$ be a class of functions parameterized by $\theta\in\Theta$ where $\Theta$ is the vector space of our parameters. Any homeomorphism $p:\Theta\to\Theta$ is a local parameter symmetry at $\theta$ if $f_{\theta}=f_{p\theta}$ and we denote the group of such transformations as $P_{\theta}$. If $p$ is a local parameter symmetry for all $\theta\in\Theta$, then we say $p$ is a global parameter symmetry and we denote the group of such transformations as $P$.
@@
For us, we usually consider $\Theta$ as our vector space and subgroups of the global parameter symmetries $H\leq P$ with a linear action. Most commonly considered parameter symmetries act linearly on parameter space.
@@
Let $f:M\to\mathbb{R}$ be a smooth function invariant under $H$. If $\Sigma$ is a submanifold, then critical points of $f|_{\Sigma}$ are critical points of $f$.
We see that the loss landscape is invariant under $P$ and in particular, invariant under any subgroup of $P$. For example, if we consider the subgroup $H$ corresponding to permuting two specific neurons, the corresponding fixed point subspace $\Theta_{H}$ corresponds to shared input and output weights for those two neurons. This is effectively a network with one fewer neuron. The principle of symmetric criticality then tells us minima of $L|_{\Theta_{H}}$ is a critical point of $L$, reminiscent of previously known results relating minima of smaller networks to critical points of larger ones (fukumizu2000local, ; simsek2021geometry, ; brea2019weight, ). To connect directly back to these prior works, we note that any transformation of the output weights of these two neurons which leaves the sum unchanged is a local symmetry of $\Theta_{H}$. Hence taking the orbit of $\Theta_{H}$ gives a larger subspace corresponding to the different ways we may embed the smaller network in a larger one.
@@ 3 Loss landscape of equivariant networks
Here we present the main theoretical results and intuition for why they are true. Proofs can be found in Appendix D. We first make some general statements that apply to general linearly constrained models, not just equivariant ones.
@@ 3.1 Hidden symmetries in constrained models
The lemma requires an orthogonality condition $(\Theta_{H}^{\perp}\cap\tilde{\Theta})+(\Theta_{H}\cap\tilde{\Theta})=\tilde{\Theta}$. This condition holds in the case of equivariant networks when we reduce a transitive p-rep to a smaller transitive p-rep and is a relatively mild condition.
While not explored in this work, we suspect Lemma 3.2 can be helpful for understanding scaling behavior of equivariant networks using loss landscape geometry arguments similar to simsek2021geometry .
If the hidden fixed point subspace $\tilde{\Theta}_{H}$ is one dimension lower than the constrained subspace, we have an even stronger statement. In this case, $\tilde{\Theta}_{H}$ in some sense splits $\tilde{\Theta}$ into two halves. We cannot traverse between the two halves via gradient descent because once we reach $\tilde{\Theta}_{H}$ we must get stuck. Generically, we expect the global minima to be in one of these halves so if we are initialized in the "bad" half, then we would never reach the global minima. This is depicted in Figure 3. Hence, we expect there to be an additional minimum on the other side.
To formally state the theorem, we require the following conditions and discuss their satisfiability.
Conditions:
  1.  C1
    $(\Theta_{H}^{\perp}\cap\tilde{\Theta})+(\Theta_{H}\cap\tilde{\Theta})=\tilde{\Theta}$
    This is used to invoke Lemma 3.2. This condition can hold in the case of equivariant networks when neuron permutation symmetries can reduce a transitive p-rep to a smaller transitive p-rep.
  2.  C2
    $\dim(\tilde{\Theta}_{H})=\dim(\tilde{\Theta})-1$
    This is a strong condition. We expect this condition can sometimes hold in the first layer but will generally fail if the previous layer is wide.
  3.  C3
    There exists some $C$ such that for any direction $\hat{r}\in\tilde{\Theta}$ we have $\hat{r}\cdot\nabla L(C\hat{r})>0$
    This condition will be satisfied if we include regularization. This is mainly used to draw a region with a minimum in its interior.
  4.  C4
    The minimum of $L|_{\tilde{\Theta}_{H}}$ is nondegenerate in $L|_{\tilde{\Theta}}$
    This is primarily used to ensure we can fully characterize saddle points and minima via the Hessian. When we have degeneracy (singular Hessian matrix), it may result from some additional symmetry which we can exploit to preserve the theorem.
@@
Suppose there is a minima $\theta_{1}$ of $L|_{\tilde{\Theta}}$ such that $\theta_{1}\notin\tilde{\Theta}_{H}$. If the conditions above hold, there must exist a distinct minimum $\theta_{2}\neq\theta_{1}$ of $L|_{\tilde{\Theta}}$.
@@ 3.2 Hidden fixed point spaces of equivariant networks
Let a $m$-neuron point be a transitive permutation representation. Suppose the weights satisfy the corresponding equivariance constraints. If this $m$-neuron point is reducible, then either we can eliminate these neurons or we can replace the transitive permutation representation with a $n$-neuron transitive permutation representation where $n<m$ and $n|m$.
@@
Let there be transitive p-reps defined on distinct blocks of $m$-neurons and $m^{\prime}$-neurons in layer $\ell$, each of which is irreducible. Suppose the $m+m^{\prime}$-neuron point combining these neurons is reducible. Then they must both be the same permutation representation and we can replace the $m+m^{\prime}$-neuron block with a $m$-neuron point.
@@ 4 Experiments
where $\sigma$ is our nonlinearity. For the experiments here, we use ReLU. Results with other activations can be found in Appendix E. All of our experiments can be run on a laptop.
@@
4.1 Permutation equivariance
In this section, we consider permutation symmetry $S_{n}$ and our example uses $S_{3}$. In Appendix E, we also consider spherical signals used in $\mathrm{O}(3)$ and $\mathrm{SO}(3)$ equivariant networks.
@@ 4.2 Loss landscape visualization
For a $\pi_{n}$ block in the hidden layer, we note there is a hidden symmetry corresponding to permuting the corresponding $n$ neurons. The resulting hidden fixed point subspace happens exactly the weights to these neurons are all the same. If we have a single $\pi_{n}$ block in the input, then the linear map $\theta_{1}I+\theta_{2}(\mathbf{1}\mathbf{1}^{T}-I)$ is in the hidden fixed point subspace when $\theta_{1}=\theta_{2}$. In particular, the degrees of freedom is reduced exactly by 1 so Theorem 3.3 applies and we expect there to be an additional induced minima. Note one can check that additional $\pi_{0}$ representations in the input or hidden layer do not affect this.
Hence, we choose $S_{3}$ and a $5$-dimensional input consisting of one $\pi_{3}$ rep and 2 trivial $\pi_{0}$ reps. In the hidden layer we choose one $\pi_{0}$ rep and $3$ trivial $\pi_{0}$ reps. We randomly generate weights for the teacher network. We then set all other weights of the student network to equal that of the teacher network and only vary the diagonal and off diagonal weights $\theta_{1}$ and $\theta_{2}$ controlling the mapping from $\pi_{3}$ to $\pi_{3}$. Figure 5 shows the resulting loss landscape for a random seed. We clearly see two distinct minima in the loss landscape. In Appendix E, we initialize teacher weights using different random seeds and see similar results.
@@ 4.3 Training, bad initializations, and parameterization
For each value of $\theta_{1},\theta_{2}$, we trained the student network for 100 steps at a learning rate of $10^{-1}$ using gradient descent on the exact loss. This is usually sufficient for the student network to converge. The resulting final loss is shown in Figure 6a. Note that indeed we see two regions, one which converges to a global minima and one which does not. However, the boundary is not a straight line. It turns out this is because gradient descent behavior depends on parameterization and in particular, on the relative scaling of the parameters (kristiadi2023geometry, ). By rescaling our parameterization to $\sqrt{n-1}\theta_{1}I+\theta_{2}(\mathbf{1}\mathbf{1}^{T}-I)$ and then performing training, we obtain Figure 6b. Here, the boundary between the two regions is indeed a straight line and corresponds to the hidden fixed point space. We emphasize, however, that no matter how we rescale our parameters, there will always be a boundary. In Appendix E, we generate teacher weights using different random seeds and see similar results.
@@ 4.4 Relaxation
Figure 8: Training of unconstrained network initialized at bad minima. We break out of the original equivariant subspace into a symmetrically related equivariant subspace. In particular, the "bad" half of the original constrained subspace is closer to the "good" half of the symmetrically related equivariant subspace. (a) Depicts values of various quantities during training of the relaxed model. (b) Projection of the parameters onto a $3$-dimensional subspace. The planes correspond to various symmetrically related equivariant subspaces color coded by good and bad halves containing the good or bad minimum respectively.
@@ 5 Conclusion
Acknowledgments and Disclosure of Funding
We acknowledge the support of the National Science Foundation under Cooperative Agreement PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions). YuQing Xie was also supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE-1745302.
We would also like to thank our helpful discussions with Robin Walters, Elyssa Hofgard, Hannah Lawrence, Vasco Portilheiro, and Yuxuan Chen.
Appendix A Notation
For convenience, we provide a table of some of the commonly used symbols in our paper and their typical meaning.
Table 1: Commonly used symbols
$G$ Group our network is equivariant under
$X$ Typically used to represent the space of our inputs
$Y$ Typically used to represent the space of our outputs
$\rho$ A group representation $\rho:G\to\mathrm{GL}(V)$
$x$ Input
$y$ Output
$\Theta$ Weight space of unconstrained networks
$f_{\theta}$ Function $f_{\theta}:X\to Y$ parameterized by $\theta\in\Theta$
$P$ Group of parameter symmetries
$\tilde{\Theta}$ Linear subspace of parameters corresponding to enforcing some constraint
$\tilde{P}$ Parameter symmetries of constrained class of networks, in particular we must have $p\tilde{\Theta}\in\tilde{\Theta}$ for all $p\in\tilde{P}$
$H$ Hidden parameter symmetry group, $H\leq P$ and $H\nleq\tilde{P}$ which arises because our choice of constraint enforcement breaks the unconstrained parameter symmetries
$\Theta_{H}$ Fixed point subspace, we have $h\theta_{H}=\theta_{H}$ for all $\theta_{H}\in\Theta_{H}$ and $g\in H$
$\tilde{\Theta}_{H}$ Hidden fixed point subspace, equal to $\Theta_{H}\cap\tilde{\Theta}_{H}$
II Identity matrix
$\mathbf{1}$ Constant matrix of 1s
$\rho$ Group representation.
$\pi$ Permutation representation.
Appendix B Other related works
Previously the optimization dynamics of equivariant, non-equivariant, and non-equivariant augmented networks were studied in [39]. This work showed that the sets of critical points of an equivariant network remain critical points when we expand to an unconstrained space if we also include data augmentation. In particular, under full augmentation, the chosen equivariant subspace is invariant under gradient flow. Our work is complementary in that we show how critical points can arise in the equivariant subspace due to interactions with parameter space symmetries. In addition, our experiments show a network leaving an equivariant subspace. This is consistent with [39] since we trained using SGD where our individual batches are not fully augmented. This means we can have gradient updates which leave the equivariant subspace.
Another very interesting series of works studies families of minima in 2-layer teacher-student ReLU setups [3, 4]. This is a similar setup to that used in our toy experiments with a model of the form
$f(x;W_{1},W_{2})=W_{2}\sigma(W_{1}x)$
where $W_{1},W_{2}$ are trainable weights and $\sigma$ is a pointwise nonlinearity chosen to be ReLU. In our setup, we fix $W_{2}$ to be a constant matrix consisting only of 1’s. In [3, 4], they choose a teacher network such that $W_{1}=I$ and $W_{2}$ consists of all 1’s. The corresponding critical points and Hessian spectrum for the loss of the student network can be analyzed by exploiting symmetries of the teacher network from which they provide a characterization of various families of minima. It turns out this analysis has intimate connections with our main theorems and the results of [39]. We briefly provide a sketch of this connection below.
In particular [4] looks at $W_{1}$ isotropic under $\Delta S_{m}\times\Delta S_{n-m}$: consisting of simultaneously permuting the first $m$ rows and columns, or simultaneously permuting the last $n=m$ rows and columns. It turns out such $W_{1}$ correspond exactly to those satisfying equivariance constraints for $G=S_{m}\times S_{n-m}$ where we choose $\pi_{n}\times\pi_{0}$ and $\pi_{0}\times\pi_{n-m}$ representations for the first $m$ and next nmn-m neurons in the hidden layer. The minima families in [4] correspond exactly to those implied by our Theorem 3.3 if constrained to these symmetric $W_{1}$; specifically weights which have blocks where diagonal values are less than the off diagonal ones. Because the input distribution is unit Gaussian and hence also symmetric under $G$, the results of [39] implies that even if we relax the constraints on $W_{1}$, the induced minima implied by our Theorem 3.3 remain critical points. The more detailed analysis of [4] show that these are minima for wide enough inputs.
Appendix C Group theory background